OpenStudy (anonymous):

Help factor!!! (x-1+i√3)(x-1-i√3)

OpenStudy (mathstudent55):

This is already factored. Do you need to multiply together?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

I know how to use the foil method with 2 terms, but not with 3 like this

OpenStudy (jdoe0001):

$$ (x-1+\sqrt{3}i)(x-1-\sqrt{3}i)\\ \text{keep in mind that => }(a-b)(a+b) = (a^2-b^2)\\ (\color{red}{x-1}+\color{blue}{\sqrt{3}i})(\color{red}{x-1}-\color{blue}{\sqrt{3}i}) $$

OpenStudy (anonymous):

I need it all multiplied together.

OpenStudy (anonymous):

Not like that lol

OpenStudy (phi):

with this, it is easiest to think of the real part and the imaginary parts as "a" and "b" in (a+b)(a-b)

OpenStudy (mathstudent55):

You can try multiplying every term of the left factor by every term of the right factor. Then collect like terms and simplify. An easier way will be if you notice that this is the product of a sum and a difference by rewriting it as: \( [(x -1) + i \sqrt{3}) ] ~[(x - 1) - i\sqrt{3} ) ] \)

OpenStudy (anonymous):

Ok one second

OpenStudy (mathstudent55):

The product of a sum and a difference. \( (a + b)(a - b) = a^2 - b^2 \) \( [(x -1) + i \sqrt{3}) ] ~[(x - 1) - i\sqrt{3} ) ]\) \(= [(x -1)^2 - (i \sqrt{3})^2 ] \)

OpenStudy (anonymous):

I got \[x ^{2}-x-ix \sqrt{3}-x+1+i \sqrt{3}+ix \sqrt{3}-i \sqrt{3}-i ^{2}\sqrt{9}\] all multiplied out. Now I need to simplify I guess

OpenStudy (phi):

yes, remember i^2 = -1

OpenStudy (mathstudent55):

Good so far.

OpenStudy (phi):

but if you look at mathstudent's post, you could do it this way (x-1)^2 - i^2 (sqrt(3))^2 the i^2 is -1 and sqrt(3) squared is 3 so (x-1)^2 + 3 which can now be expanded (relatively) easily x^2 -2x+1 + 3 x^2 -2x +4

OpenStudy (anonymous):

Well, I'm down to \[x ^{2}-2x+2\sqrt{9}\]

OpenStudy (anonymous):

The square root of 9 is 3, so would that be \[x ^{2}-2x+2*3\] ?

OpenStudy (anonymous):

To get the final answer of \[x ^{2}-2x+6\]?

OpenStudy (phi):

you should have gotten down to \[x^2 -2x + 1 + \sqrt{9} \] double check what you did

OpenStudy (anonymous):

1-\[i ^{2}\]= 1-(-1) = 1+1 = 2

OpenStudy (anonymous):

I'm pretty sure it is 2

OpenStudy (phi):

yes but it is 1 - i^2 * sqrt(9) which is 1 - -1*sqrt(9) = 1 + sqrt(9)

OpenStudy (anonymous):

It's a double negative so it turns to a positive

OpenStudy (mathstudent55):

\((x-1+i\sqrt{3})(x-1-i\sqrt{3}) \) \(x ^{2}-x-ix \sqrt{3}-x+1+i \sqrt{3}+ix \sqrt{3}-i \sqrt{3}-i ^{2}\sqrt{9}\) \(x^2 - x - x -ix \sqrt{3} + ix \sqrt{3} + 1 + 3\)

OpenStudy (phi):

yes but it is not 1 - i^2 it is \[ 1 - i^2 \sqrt{9} \]

OpenStudy (anonymous):

Which is \[1-(-1)\sqrt{9}\] lol

OpenStudy (phi):

order of operations means you do i^2 to get -1 then multiply -1* sqrt(9) to get -sqrt(9) now you have 1 - -sqrt(9) which becomes 1 + sqrt(9)

OpenStudy (anonymous):

Ohhhhhh ok ok. Wasn't thinking in terms of orders of operations for some reason haha

OpenStudy (anonymous):

So the final answer would be \[x ^{2}-2x+4\]

OpenStudy (phi):

yes. but if you go back to the previous posts you will see the fastest way to do this is treat the real part (x+1) and the imaginary parts as (a+b i) (a- bi) these are *complex conjugates* and you get a^2 - i^2 b^2 which is a^2 + b^2 try to memorize that: complex conjugates give you the sum of the squares (pure real) so (x+1)^2 + 3 which becomes x^2 -2x +4

OpenStudy (anonymous):

Yeah, that would have saved me a lot of time. Thanks so much

OpenStudy (phi):

it is worth doing it the hard way... it motivates learning the short cut.

OpenStudy (anonymous):

So true! Have a great rest of the day